Kingston Company uses the dollar-value LIFO method of computing inventory. An external price index is used to convert ending inventory to a base year. The company began operations on January 1, 2021, with an inventory of $165,000. Year-end inventories at year-end costs and cost indexes for its one inventory pool were as follows:

Year Ended Ending Inventory Cost Index
December 31 at Year-End Costs (Relative to Base Year)
2021 $ 243,800 1.06
2022 324,500 1.18
2023 304,750 1.15
2024 299,700 1.11


Required:
Calculate inventory amounts at the end of each year. (Round intermediate calculations and final answers to the nearest whole dollars.)
Answer #1

Calculate ending inventory

Ending inventory
1/1/2021 165000
12/31/2021 243800/1.06 = 230000 (165000+65000*1.06) = 233900
12/31/2022 324500/1.18 = 275000 (233900+45000*1.18) = 287000
12/31/2023 304750/1.15 = 265000 (233900+35000*1.18) = 275200
12/31/2024 299700/1.11 = 270000 (275200+5000*1.11) = 280750